detemine the value of expresion
5^3-27=
7^77-7^76=

Answer by stanbon(75887)

detemine the value of expresion
5^3-27= 125-27 = 98
7^77-7^76= 7^76(1-7)=6(7^76)= 1.01298 x 10^65
Cheers,
Stan H.